OpenStudy (smarty):

A rocket is shot straight into the air at an initial velocity of 120 feet/second. The height of the rocket after t seconds is given by the equation h = 120t – 4t2. After how many seconds will the rocket reach its maximum height?

OpenStudy (anonymous):

Vertex of a parabola: -b/2a

OpenStudy (anonymous):

first you must rearrange your equation in receding powers

OpenStudy (smarty):

A.900 seconds B.120 seconds C.30 seconds D.15 seconds

OpenStudy (anonymous):

This can be solved using calculus. Given h = 120 t - 4 t^2 Differentiate, dh/dt = 120 - 8t Differentiate a 2nd time, d2h/dt2 = -8 This shows that whatever stationary point it is a maximum turning point, which is what you want. Returning to dh/dt equate dh/dt= 0 to find the value of t which makes h maximum. 120-8t = 0 8t=120 t=15 When t = 15, h is maximum

OpenStudy (smarty):

Thanks, so D?

OpenStudy (anonymous):

@tiaph , first of all you gave the answer without any talk about how to go about the problem. This caused you to do it in a calculus fashion when there is two ways to solve it. No where did you help smarty come to this answer as you never asked what class he was taking.

OpenStudy (anonymous):

Because i saw you typing "vertex of the parabola... " so i assume you are gonna offer the geometrical or graphing method to solve the question. I am just using calculus to provide a 2nd method of answering. I assure you when i answer question i do my best to make my explanation as clear and succinct as possible, so people understand what i am doing. I apologize, if my seemingly rash and unconcerned answer provide much distress to you.